Sketch the graph of f(x)=|x| where x belongs to R, state it domain and range.


1
Expert's answer
2022-04-18T16:51:00-0400

"Dom\left( f \right) =\mathbb{R} \\Range\left( f \right) =\left[ 0,+\infty \right)"